# Misspecification-robust amortised simulation-based inference using variational methods

### v2

@@ 1 Introduction
A range of SBI methods have emerged to solve the likelihood-free inference problem, beginning with traditional approaches such as approximate Bayesian computation (ABC; 10.1214/aos/1176346785; abc_genetics) and Bayesian synthetic likelihood (BSL, Price02012018). Recent work has introduced methods based on neural density estimation, such as neural posterior estimation (NPE, 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), neural likelihood estimation (NLE, pmlr-v96-lueckmann19a; papamakarios2019sequentialneurallikelihoodfast), neural ratio estimation (NRE, pmlr-v33-izbicki14; cranmer2016approximatinglikelihoodratioscalibrated; hermans2020likelihoodfreemcmcamortisedapproximate; pmlr-v119-durkan20a), and modelling the joint distribution of data and simulation parameters (DBLP:conf/icml/GlocklerDWWM24). SBI methods can be categorised into amortised and non-amortised inference methods. In the context of neural SBI, non-amortised methods such as sequential neural posterior estimation (SNPE; 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), sequential neural likelihood estimation (SNLE; papamakarios2019sequentialneurallikelihoodfast), and sequential neural ratio estimation (SNRE; hermans2020likelihoodfreemcmcamortisedapproximate) target a single posterior conditioned on fixed data, adapting their inference procedure with each simulation round. After an up-front simulation budget, amortised methods aim to learn a global inference model over a given prior, making them well-suited for scenarios where repeated or scalable inference is required. In this paper, we focus on SBI methods for inferring posterior distributions that are amortised over a dataset of observations.
@@ 1.1 Our contributions
Amortised SBI: In this paper, we refer to amortised SBI as amortising the Bayesian inference over all data points in $\mathcal{O}$. This definition is in line with amortised variational inference (margossian2023amortised).
Our solution: We propose robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T), methods that address the misspecification problem in amortised SBI. RVNP pre-trains a simulator likelihood $p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$, adopts a flexible error model $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, and uses an importance weighted autoencoder (burda2015importance) inference scheme to maximize the evidence of the true data under the variational posterior $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ for the parameters $\bm{\xi}$ and $\bm{\phi}$.
@@ 2.1 SBI formalism
We define the error model as $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, a family of distributions parametrized by $\bm{\xi}(\bm{\theta})\in\mathbb{R}^{k}$ and $\bm{x}_{\rm sim}$, which relates the simulations to observations. The density explicitly depends on $\bm{x}_{\rm sim}$ and implicitly on $\bm{\theta}$ due to the error model parameters $\bm{\xi}$ being a function of $\bm{\theta}$.
@@ 2.2 Amortised neural posterior and neural likelihood estimation
$\mathcal{L}_{\rm NPE}({\bm{\phi},\bm{\omega}})$ $=\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{KL}[p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))||q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big]$ (1)
$=\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{E}_{p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))}[-\log q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big],$
where the expectation is over the unknown true data-generating distribution $p^{*}(\bm{x}_{\rm obs})$. The second line follows from the first because the true unknown posterior and its entropy do not depend on the trainable parameters.
@@
$\mathcal{L}(\bm{\Psi})_{\rm NLE}:=\mathbb{E}_{p(\bm{x}_{\rm sim}|\bm{\theta})p(\bm{\theta})}[\log p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\bm{\theta})]$ (3)
with respect to the flow parameters $\bm{\Psi}$ (where we dropped reference to $\iota_{\bm{\omega}}$), which is equivalent to minimising the KL divergence between the flow and the target distribution. Using neural embeddings for the summary statistics must be implemented in a two-stage manner because it cannot be trained in an end-to-end fashion along with the likelihood proxy (brehmer2020flows). Sampling from the posterior amounts to adopting either a Markov Chain Monte Carlo (MCMC) sampling method or a variational inference algorithm, which will increase the number of architecture choices and hyperparameter tuning, but can recover accurate posterior inference (glockler2022variational).
@@ 2.3 Misspecification in SBI
The simulator, $p(\bm{x}_{\rm sim}|\,\bm{\theta})$, is said to be misspecified if the true data generative process does not fall within the family of distributions defined by the simulator on the support of the prior of $\bm{\theta}$. That is, $q^{*}\notin\{p(x_{\rm sim}|\,\bm{\theta})\,;\bm{\theta}\in\bm{\Theta})\}$ (cannon2022investigating). For single observations, the definition of misspecification can be extended to using summary statistics (kelly2024misspecificationrobust) by defining $\bm{b}(\bm{\theta})=\mathbb{E}_{p(\bm{x}_{\rm sim}\mid\bm{\theta})}[\iota_{\omega}(\bm{x}_{\rm sim})]$ and $\bm{b}_{j}=\mathbb{E}_{q^{*}(\bm{x}_{\rm obs})}[\iota_{\omega}(\bm{x}_{\rm obs}^{(j)})]$, for each $\bm{x}_{\rm obs}^{(j)}\in O$ where $O=\{\bm{x}_{\rm obs}^{(i)}\}_{i=1}^{N_{\rm obs}}$ is the set of observations each associated to a different, unknown true $\bm{\theta^{(j)}}$ value. The simulator is said to be misspecified if there is no $\{\bm{\theta}\}_{i=1}^{N_{\rm obs}}$ in the support of the prior for which $\bm{b}(\bm{\theta}^{(j)})=\bm{b}_{j}$ for each $j$.
In this paper, we adopt the alternate, but similar, definition of misspecification provided in wehenkel2025addressing, which is more aligned with amortised SBI. They define a simulator to be misspecified if $\exists\mathcal{S}\subseteq\Theta\times\mathcal{X}_{\rm obs}:\forall(\bm{\theta},\bm{x}_{\rm obs})\in\mathcal{S}$
@@ 2.4 Importance weighted autoencoders
Importance-weighted autoencoders (IWAE) were inspired by the vanilla variational autoencoder (VAE, kingma2013auto), which introduces a latent space ${\Theta}\subseteq\mathbb{R}^{m}$ and two distributions, $p_{\bm{\xi}}(\bm{x}|\,\bm{\theta})$ and $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x})$, known as the decoder and the encoder respectively. The vanilla VAE aims to maximise the evidence of the data using the evidence lower bound
$\log p(\mathbf{x})\;\geq\;\mathbb{E}_{q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})}\left[\log p_{\bm{\xi}}(\mathbf{x}|\mathbf{\theta})\right]-\mathrm{KL}\!\left(q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})\,\|\,p(\mathbf{\theta})\right)$ (5)
with respect to $\bm{\xi}$ and $\bm{\phi}$, where $p(\bm{\theta})$ is a prior over the latent parameter. However, using this bound has been shown to induce mode-seeking behaviour, leading to overly simplified representations and poor inference. burda2015importance introduced IWAE, which is based on a strictly tighter evidence lower bound derived from importance sampling. The log-evidence for a single point is given as
$\log p_{\bm{\theta}}(\bm{x})$ $\approx\log\left[\frac{1}{k}\sum_{l=1}^{k}\frac{p_{\bm{\xi}}\left(\textbf{x},\bm{\theta}^{(l)}\right)}{q_{\bm{\phi}}\left(\bm{\theta}^{(l)}|\textbf{x}\right)}\right]={\mathcal{L}}^{\text{IWAE}}_{k}\left(\bm{\xi},\bm{\phi};\textbf{x}\right)$ (6)
@@ 3.1 Generative model
Throughout, we assume a pre-training or fixing of the neural statistic embedding $\iota_{\omega}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$ occurred. The first step of RVNP is to train the normalising flow $p_{\bm{\Psi}}(\bm{x}|\,\bm{\theta})$ to approximate the likelihood using the generated samples from the simulator by minimising the NLE objective (Equation 3). Once the normalising flow has been trained, we assume that $p(\bm{x}_{\rm sim}|\,\bm{\theta})\approx p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$ and include the surrogate in the forward model.
We assume that the true DGP can be modelled as
@@
    $p(\bm{\xi})$ a (pseudo-)prior over the error model parameters.
@@
    $\bm{x}_{\rm obs}\sim p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, where $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\mathcal{N}(\bm{x}_{\rm obs};\,\bm{x}_{\rm sim},\bm{\xi}(\bm{\theta}))$ is an adopted error model conditional on $\bm{\xi}$, where the covariance matrix is the output of a neural network $\rm{NN}(\bm{\theta})$ parametrized by $\bm{\alpha}$.
Under this generative model, the posterior distribution is proportional to
$p(\{\bm{\theta}^{(i)}\}_{i=1}^{N_{\rm obs}},\bm{\xi}\mid O)\propto\prod_{i=1}^{N_{\rm obs}}\int p_{\bm{\xi}(\bm{\theta}^{(i)})}(\bm{x}^{(i)}_{\rm obs}\mid\bm{x}^{(i)}_{\rm sim})\,p_{\bm{\Psi}}(\bm{x}^{(i)}_{\rm sim}\mid\bm{\theta}^{(i)})\,p(\bm{\theta}^{(i)})\,p(\bm{\xi}(\bm{\theta}^{(i)}))\,d\bm{x}^{(i)}_{\rm sim}$. (10)
@@ 3.2 Variational posterior
$\log p(O)=\sum_{i=1}^{N_{\rm obs}}\log\mathbb{E}_{p_{\bm{\Psi}}(\bm{x}^{(i)}_{\rm sim}|\,\bm{\theta}^{(i)})}[p_{\bm{\xi}(\bm{\theta}^{(i)})}(\bm{x}^{(i)}_{\rm obs}|\,\bm{x}^{(i)}_{\rm sim})p(\bm{\theta}^{(i)})p(\bm{\xi}(\bm{\theta}^{(i)}))].$ (11)
@@
$\mathcal{L}(\bm{\phi},\bm{\alpha})_{V}\approx-\sum_{i=1}^{N_{\rm obs}}\log\left[\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]\,p(\bm{\theta}^{(l)})p(\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha}))}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}\right]$ (12)
where $\bm{\theta}^{(l)}\sim q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)})$, $\bm{\alpha}$ are the weights of the neural network $\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})$, and for each $\bm{\theta}^{(l)}$ we approximate $\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]$ using a Monte Carlo (MC) estimate. We use the logsumexp function to ensure the MC estimates are stable.
@@ 3.3 Posterior tuning
$\mathcal{L}(\bm{\phi})_{\rm{NPE}(\alpha)}:=-\mathbb{E}_{p(\bm{x}_{\rm sim},\bm{\theta})}\mathbb{E}_{p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})}[\log q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})].$ (13)

### v3

@@ 1 Introduction
A range of SBI methods have emerged to solve the likelihood-free inference problem, beginning with traditional approaches such as approximate Bayesian computation (ABC; 10.1214/aos/1176346785; abc_genetics) and Bayesian synthetic likelihood (BSL, Price02012018). Recent work has introduced methods based on neural density estimation, such as neural posterior estimation (NPE, 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), neural likelihood estimation (NLE, pmlr-v96-lueckmann19a; papamakarios2019sequentialneurallikelihoodfast), neural ratio estimation (NRE, pmlr-v33-izbicki14; cranmer2016approximatinglikelihoodratioscalibrated; hermans2020likelihoodfreemcmcamortizedapproximate; pmlr-v119-durkan20a), and modelling the joint distribution of data and simulation parameters (DBLP:conf/icml/GlocklerDWWM24). SBI methods can be categorised into amortised and non-amortised inference methods. In the context of neural SBI, non-amortised methods such as sequential neural posterior estimation (SNPE; 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), sequential neural likelihood estimation (SNLE; papamakarios2019sequentialneurallikelihoodfast), and sequential neural ratio estimation (SNRE; hermans2020likelihoodfreemcmcamortizedapproximate) target a single posterior conditioned on fixed data, adapting their inference procedure with each simulation round. After an up-front simulation budget, amortised methods aim to learn a global inference model over a given prior, making them well-suited for scenarios where repeated or scalable inference is required. In this paper, we focus on SBI methods for inferring posterior distributions that are amortised over a dataset of observations.
@@ 1.1 Our contributions
Amortised SBI: In this paper, we refer to amortised SBI as amortising the Bayesian inference over all data points in $\mathcal{O}$. This definition is in line with amortised variational inference (margossian2023amortized).
Our solution: We propose robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T), methods that address the misspecification problem in amortized SBI. RVNP pre-trains a simulator likelihood $p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$, adopts a flexible error model $p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})$, and uses an importance weighted autoencoder (burda2015importance) inference scheme to maximize the evidence of the true data under the variational posterior $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ for the parameters $\bm{\alpha}$ and $\bm{\phi}$.
@@ 2.1 SBI formalism
We define the error model as $p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})$, a family of distributions parametrized by $\bm{\alpha}\in\mathbb{R}^{k}$, $\bm{x}_{\rm sim}$, and $\bm{\theta}$, which relates the simulations to observations.
@@ 2.2 Amortised neural posterior and neural likelihood estimation
$\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{KL}[p(\bm{\theta}|\,\bm{x}_{\rm obs})||q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big],$ (1)
where the expectation is over the unknown true data-generating distribution $p^{*}(\bm{x}_{\rm obs})$. Minimizing this is equivalent to minimizing the NPE loss function
$\mathcal{L}_{\rm NPE}({\bm{\phi},\bm{\omega}})$ $=\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{E}_{p(\bm{\theta}|\,\bm{x}_{\rm obs})}[-\log q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})]\big],$ (2)
since the true unknown posterior and its entropy do not depend on the trainable neural network parameters.
@@
$\mathcal{L}(\bm{\Psi})_{\rm NLE}:=\mathbb{E}_{p(\bm{x}_{\rm sim}|\bm{\theta})p(\bm{\theta})}[\log p_{\bm{\Psi}}(\iota_{\bm{\omega}}(\bm{x}_{\rm sim})|\bm{\theta})]$ (4)
with respect to the flow parameters $\bm{\Psi}$, which is equivalent to minimising the KL divergence between the flow and the target distribution. The parameters for the neural summary statistics must be implemented in a two-stage manner because it cannot be trained in an end-to-end fashion along with the likelihood proxy (brehmer2020flows). Sampling from the posterior amounts to adopting either a Markov Chain Monte Carlo (MCMC) sampling method or a variational inference algorithm, which will increase the number of architecture choices and hyperparameter tuning, but can recover accurate posterior inference (glockler2022variational).
@@ 2.3 Misspecification in SBI
The simulator, $p(\bm{x}_{\rm sim}|\,\bm{\theta})$, is said to be misspecified if the true data generative process does not fall within the family of distributions defined by the simulator on the support of the prior of $\bm{\theta}$. That is, $p^{*}\notin\{p(x_{\rm sim}|\,\bm{\theta})\,;\bm{\theta}\in\bm{\Theta})\}$ (cannon2022investigating). In this paper, we adopt the adjusted definition of misspecification provided in wehenkel2025addressing, which is more aligned with amortised SBI. They define a simulator to be misspecified if $\exists\mathcal{S}\subseteq\Theta\times\mathcal{X}_{\rm obs}:\forall(\bm{\theta},\bm{x}_{\rm obs})\in\mathcal{S}$
@@ 2.4 Importance weighted autoencoders
Importance-weighted autoencoders (IWAE) were inspired by the vanilla variational autoencoder (VAE, kingma2013auto), which introduces a latent space ${\Theta}\subseteq\mathbb{R}^{m}$ and two distributions, $p_{\bm{\alpha}}(\bm{x}|\,\bm{\theta})$ and $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x})$, known as the decoder and the encoder respectively. The vanilla VAE aims to maximise the evidence of the data using the evidence lower bound
$\log p(\mathbf{x})\;\geq\;\mathbb{E}_{q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})}\left[\log p_{\bm{\alpha}}(\mathbf{x}|\mathbf{\theta})\right]-\mathrm{KL}\!\left(q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})\,\|\,p(\mathbf{\theta})\right)$ (6)
with respect to $\bm{\alpha}$ and $\bm{\phi}$, where $p(\bm{\theta})$ is a prior over the latent parameter. However, using this bound has been shown to induce mode-seeking behaviour, leading to overly simplified representations and poor inference. burda2015importance introduced IWAE, which is based on a strictly tighter evidence lower bound derived from importance sampling. The log-evidence for a single point is given as
$\log p_{\bm{\theta}}(\bm{x})$ $\approx\log\left[\frac{1}{k}\sum_{l=1}^{k}\frac{p_{\bm{\alpha}}\left(\textbf{x},\bm{\theta}^{(l)}\right)}{q_{\bm{\phi}}\left(\bm{\theta}^{(l)}|\textbf{x}\right)}\right]={\mathcal{L}}^{\text{IWAE}}_{k}\left(\bm{\alpha},\bm{\phi};\textbf{x}\right)$ (7)
@@ 3.1 Generative model
Throughout, we assume a pre-training or fixing of the neural statistic embedding $\iota_{\omega}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$ occurred. The first step of RVNP is to train the normalising flow $p_{\bm{\Psi}}(\bm{x}|\,\bm{\theta})$ to approximate the likelihood using the generated samples from the simulator by minimising the NLE objective (Equation 4). Once the normalising flow has been trained, we assume that $p(\bm{x}_{\rm sim}|\,\bm{\theta})\approx p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$ and include the surrogate in the forward model. We assume that the true DGP can be modelled as
@@
    $p(\bm{\alpha})$ a prior over the error model weights.
@@
    $\bm{x}_{\rm obs}\sim p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})$, where $p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})=\mathcal{N}(\bm{x}_{\rm obs};\,\bm{x}_{\rm sim},\bm{\xi}(\bm{\theta}\,;\,\bm{\alpha}))$ is an adopted error model where the covariance matrix is the output of a neural network $\rm{NN}(\bm{\theta})$ parametrized by $\bm{\alpha}$.
We assume that the observed data $\bm{x}_{\rm obs}^{(i)}$ are conditionally independent given their latent parameters $\bm{\theta}^{(i)}$ and the error model parameters $\bm{\alpha}$. Under this generative model, the posterior distribution is proportional to
$p(\{\bm{\theta}^{(i)}\}_{i=1}^{N_{\rm obs}},\bm{\alpha}\mid O)\propto p(\bm{\alpha})\prod_{i=1}^{N_{\rm obs}}\int p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})\,p_{\bm{\Psi}}(\bm{x}^{(i)}_{\rm sim}\mid\bm{\theta}^{(i)})\,p(\bm{\theta}^{(i)})\,d\bm{x}^{(i)}_{\rm sim}$. (11)
@@ 3.2 Variational posterior
$\log p(O)=\sum_{i=1}^{N_{\rm obs}}\log\mathbb{E}_{p(\bm{\theta}^{(i)})}\mathbb{E}_{p_{\bm{\Psi}}(\bm{x}^{(i)}_{\rm sim}|\,\bm{\theta}^{(i)})}[p_{\bm{\alpha}}(\bm{x}^{(i)}_{\rm obs}|\,\bm{x}^{(i)}_{\rm sim},\bm{\theta}^{(i)})]+\log p(\bm{\alpha}).$ (12)
@@
$\mathcal{L}(\bm{\phi},\bm{\alpha})_{V}\approx-\sum_{i=1}^{N_{\rm obs}}\log\left[\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\alpha}}(\bm{x}^{(i)}_{\rm obs}|\,\bm{x}^{(i)}_{\rm sim},\bm{\theta}^{(l)})\right]\,p(\bm{\theta}^{(l)})}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}\right]-\log p(\bm{\alpha})$ (13)
where $\bm{\theta}^{(l)}\sim q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)})$, and for each $\bm{\theta}^{(l)}$ we approximate $\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\alpha}}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim},\bm{\theta}^{(l)}\right)\right]$ using a Monte Carlo (MC) estimate. We use the logsumexp function to ensure the MC estimates are stable.
@@ 3.3 Posterior tuning
$\mathcal{L}(\bm{\phi})_{\rm{NPE}(\alpha)}:=-\mathbb{E}_{p(\bm{x}_{\rm sim},\bm{\theta})}\mathbb{E}_{p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})}[\log q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})].$ (14)